OpenStudy (anonymous):

sqrt{3x-5}=2-sqrt{x-1}

OpenStudy (mertsj):

Square both sides.

OpenStudy (mertsj):

\[3x-5=4-4\sqrt{x-1}+x-1\]

OpenStudy (anonymous):

find all real and imaginary solutions to equation

OpenStudy (mertsj):

\[2x-8=4\sqrt{x-1}\]

OpenStudy (mertsj):

Now divide both sides by 2 and then square both sides again.

OpenStudy (mertsj):

\[x-4=2\sqrt{x-1}\]

OpenStudy (kropot72):

\[2x-8=-4\sqrt{x-1}\]

OpenStudy (mertsj):

\[x^2-8x+16=4(x-1)\]

OpenStudy (anonymous):

mertsj did you see kropots answer it would be -4

OpenStudy (mertsj):

@kropot72 The negative makes no difference. You're about to square it anyway.

OpenStudy (anonymous):

gotcha ty

OpenStudy (mertsj):

If it makes you feel better, we can include it? \[x-4=-2\sqrt{x-1}\]

OpenStudy (mertsj):

Now square both sides.

OpenStudy (anonymous):

no i see your point

OpenStudy (anonymous):

\[x ^{2}=12x-20\]

OpenStudy (kropot72):

Collect all the terms onto the left hand side and you will have a quadratic to solve

OpenStudy (anonymous):

x-10 and x-2 are my answers

OpenStudy (kropot72):

x = 10 and x = 2 are the solution

OpenStudy (anonymous):

ty kropot72 and mertsj

OpenStudy (kropot72):

All credit to mertsj.